OpenStudy (anonymous):

B is a 4x4 matrix. det(B) = -5 what is det(B^4)?

OpenStudy (anonymous):

det(b^4)=(det(b))^4=-5^4=625

OpenStudy (anonymous):

thought so. gracias!

OpenStudy (anonymous):

No problem :P Good answer please :3

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

Thank you :D That lets us know which questions have been answered xP

OpenStudy (anonymous):

oh ok will do from now on. do you get something for answering and helping other people? or just a good conscious lol